# Automating Android Build Repair: Bridging the Reasoning-Execution Gap in LLM Agents with Domain-Specific Tools

### v1

@@ Appendix B Validation of the Gemini-CLI Replica
In our experiments, we required the ability to systematically modify the toolset available to the agent to isolate the impact of domain-specific tools. To facilitate these controlled experiments, we developed a Python replica of the open-source Gemini-CLI agent (Google Cloud, ). Our replica preserves the core agentic loop and prompting structure of the original implementation while allowing for straightforward customization of the available tools. To validate the fidelity of our replica, we conducted an experiment to compare its performance against the official, unmodified Gemini-CLI agent. For a fair comparison, both agents were configured to use only the general-purpose shell tool, in addition to standard file system tools. This configuration tests whether our replica preserves the capabilities of the original agent framework. The results of this validation are shown in Table 9. We find that the resolve rates are nearly identical. This small difference is well within the expected variance due to model stochasticity. Temperature is kept at the default of 1. The close alignment in performance confirms that our replica is an accurate and reliable reproduction of the original agent, validating its use for the Gemini-CLI (Read/Write Only) and GradleFixer configurations in our experiments.

### v2

@@ Appendix B Validation of the Gemini-CLI Replica
In our experiments, we required the ability to systematically modify the toolset available to the agent to isolate the impact of domain-specific tools. To facilitate these controlled experiments, we developed a Python replica of the open-source Gemini-CLI agent (google_gemini_cli). Our replica preserves the core agentic loop and prompting structure of the original implementation while allowing for straightforward customization of the available tools. To validate the fidelity of our replica, we conducted an experiment to compare its performance against the official, unmodified Gemini-CLI agent. For a fair comparison, both agents were configured to use only the general-purpose shell tool, in addition to standard file system tools. This configuration tests whether our replica preserves the capabilities of the original agent framework. The results of this validation are shown in Table 9. We find that the resolve rates are nearly identical. This small difference is well within the expected variance due to model stochasticity. Temperature is kept at the default of 1. The close alignment in performance confirms that our replica is an accurate and reliable reproduction of the original agent, validating its use for the Gemini-CLI (Read/Write Only) and GradleFixer configurations in our experiments.